OpenStudy (anonymous):

X^4-81/x+3 state the quotient is and remainder. I got the quotient is x^3-3x^2+9x-27 and no remainder. Please check my math. Thank you.

OpenStudy (anonymous):

you are right, there is no remainder

OpenStudy (anonymous):

\[x^4-81=(x^2-9)(x^2+9)=(x-3)(x+3)(x^2+9)\]

OpenStudy (anonymous):

Thank you^_^.